OpenStudy (anonymous):

What is the vertex of the function f(x) = 2x2 + 12x + 9? HELP!?

OpenStudy (anonymous):

\[ 2(x^2 -6x +(\frac{6}{2})^2 ) +9 - (2)(\frac{6}{2})^2 \]

OpenStudy (anonymous):

\[ = 2(x-9)^2 +18 \\ \text{Vertex is} (9,18)\]

OpenStudy (anonymous):

thanks c: